### v1

# Multimodal Large Language Models to Support Real-World Fact-Checking

Abstract
Multimodal large language models (MLLMs) carry the potential to support humans in processing vast amounts of information. While MLLMs are already being used as a fact-checking tool, their abilities and limitations in this regard are understudied. Here is aim to bridge this gap. In particular, we propose a framework for systematically assessing the capacity of current multimodal models to facilitate real-world fact-checking. Our methodology is evidence-free, leveraging only these models’ intrinsic knowledge and reasoning capabilities. By designing prompts that extract models’ predictions, explanations, and confidence levels, we delve into research questions concerning model accuracy, robustness, and reasons for failure. We empirically find that (1) GPT-4V exhibits superior performance in identifying malicious and misleading multimodal claims, with the ability to explain the unreasonable aspects and underlying motives, and (2) existing open-source models exhibit strong biases and are highly sensitive to the prompt. Our study offers insights into combating false multimodal information and building secure, trustworthy multimodal models. To the best of our knowledge, we are the first to evaluate MLLMs for real-world fact-checking.
1 Introduction
Misinformation is a significant challenge on the internet, especially with regard to multimodal claims, which combine text, images, videos, and other media types Mubashara et al. (2023). The visual component in these cases could be manipulated, used out-of-context, or accompanied by text that makes a false claim Huh et al. (2018); Luo et al. (2021); Aneja et al. (2023); Yao et al. (2023). In such cases, fact-checkers and the tools they employ need to be able to handle multiple modalities.
Large language models (LLMs) store extensive information about the world, which exceeds what any individual human can know and is accessible in a far more human-friendly way than any search engine can offer Zhao et al. (2023); Hu et al. (2023a). As such, they can be a powerful tool in the hands of fact-checkers, who often seek extra factual knowledge to verify a claim Nakov et al. (2021). The application of LLMs to fact-checking has been extensively studied in a text-only setting, wherein a model receives as input both a textual claim, with or without a set of evidence documents, and needs to reason to determine the veracity of the claim Chen et al. (2022); Pan et al. (2023); Li et al. (2023a). LLMs have increasingly strong understanding and reasoning capabilities, so using them to sift through evidence documents when verifying a claim is proving very effective Chen et al. (2022); Li et al. (2023a). Yet, this pipeline approach requires an extra step of evidence retrieval, which is computationally heavy and error-prone, while it overlooks the vast amount of knowledge stored in the parameters of LLMs Hu et al. (2023b). Additionally, LLMs are now giving way from text-only to multimodal models, which can process signals from multiple sources and perform a wider range of tasks Liu et al. (2023b); Zhang et al. (2023); Yin et al. (2023). We hypothesize that multimodal large language models (MLLMs) trained on large amounts of data can serve as a sufficient substitute for retrieving evidence and perform fact-checking in an evidence-free fashion, i.e., relying solely on their parametric knowledge.
Despite their great promise, however, MLLMs have limitations as well, specifically with respect to factuality Fu et al. (2023); Liu et al. (2023c); Xu et al. (2023). Models may produce hallucinations Huang et al. (2023); Li et al. (2023b) and suffer from a lack of knowledge or exhibit biases Hu et al. (2023a); Wang et al. (2023). To understand MLLMs’ potential to support fact-checking real-world claims, we propose the evaluation framework illustrated in Figure 1. It involves querying models for predictions, explanations, and confidence scores on multimodal check-worthy claims. Based on the data obtained, we can further assess the models’ capabilities from various perspectives, addressing the following research questions:
RQ1: How accurate are MLLMs in identifying misinformation in multimodal claims?
RQ2: How do different modalities (text and image) contribute to the assessment of multimodal content?
RQ3: Are models sensitive to prompts in different languages or expressions, or do they exhibit robustness?
RQ4: Can in-context learning enhance model performance?
RQ5: Can MLLMs accurately estimate the correctness of their responses in terms of a confidence score?
RQ6: What types of errors are most common for MLLMs employed as multimodal fact-checkers?
Our investigation spans an array of vision-language models, including GPT-4V, LLaVA-1.5 Liu et al. (2023b), MiniGPT-v2 Chen et al. (2023), and InstructBLIP Dai et al. (2023), which stand as leaders in this domain. We evaluate these models across three existing multimodal fact-checking datasets Zlatkova et al. (2019); Aneja et al. (2023); Yao et al. (2023), and one new dataset, which we build from recent check-worthy claims to mitigate the risk of data contamination.
Our study finds that MLLMs display a range of fact-checking capabilities. GPT-4V, in particular, excels in our tests, exhibiting high accuracy, useful explanations, and good calibration, with the accuracy on different datasets generally reaching around . On the other hand, open-source models like MiniGPT-v2 and InstructBLIP lack explanatory and uncertainty-reporting capabilities. LLaVA models show a strong sensitivity to prompt formulation, a tendency to comply with the question in the prompt, and hallucinatory behaviors.
In summary, we systematically evaluate existing multimodal models to determine their capability to support real-world fact-checking. We formulate several basic research questions and assess their accuracy, robustness, calibration, and reasons for failure across various datasets. Our work fills a research gap, providing a comprehensive overview of the domain and shedding light on both fact-checking and MLLMs’ factuality. As far as we are aware, our study is the first assessment of MLLMs in the context of real-world fact-checking.
2 Background and Related Work
LLMs for text-only fact-checking
Hoes et al. (2023) find that ChatGPT accurately classifies 69% of statements in a dataset built from the fact-checking platform PolitiFact. FactLLaMA Cheung and Lam (2023) integrates external evidence into the instruct-tuning process to enhance the model’s ability to leverage evidence for predictions. Studies have also focused on leveraging the capabilities of LLMs to develop automated frameworks for decomposing claims, generating queries, and predicting based on gathered evidence, thereby augmenting the interpretability of the fact-checking process Pan et al. (2023); Li et al. (2023a). However, these works often focused solely on prediction accuracy. Our work analyzes the model’s fact-checking capabilities from three perspectives: predictions, explanations, and confidence.
Multimodal fact-checking
Multimodal misinformation has been found to spread faster than its text-only counterparts Li and Xie (2020). It primarily falls into two categories: one involves textual claims about manipulated content across different modalities, while the other pertains to out-of-context misinformation, featuring unaltered images, audio, or videos in misleading contexts Luo et al. (2021); Aneja et al. (2023); Yao et al. (2023). The multimodal fact-checking is gnerally structured into distinct stages within pipelines, including claim detection extraction, evidence retrieval, verdict prediction, and explanation Nakov et al. (2021); Mubashara et al. (2023). In our work, we posit that evidence is parametrically encoded within the models, enabling us to bypass the evidence retrieval phase and directly engage in predictions and explanations for the targeted claims.
Evaluation of MLLMs
With the advancement of MLLMs, the systematic evaluation of their capabilities has become increasingly critical. A broad array of benchmarks has been established to estimate the MLLM’s performance across a variety of tasks, including perceptual tasks such as image recognition, image description, and optical character recognition, as well as cognitive tasks like commonsense reasoning and numerical calculation Yin et al. (2023); Fu et al. (2023); Liu et al. (2023c); Xu et al. (2023); Tong et al. (2024). Tu et al. (2023) access the resilience of MLLMs against various image-related attacks from a security perspective. However, there is an absence of research on evaluating the ability to conduct multimodal fact-checking with MLLMs. Our work aims to bridge this gap.
3 Evaluation Framework
We propose an evaluation framework that analyzes model capabilities along the various dimensions illustrated in Figure 1. Our experiments are conducted in an evidence-free setting, meaning that models have access only to the multimodal claims without any external knowledge. Through extensive controlled experiments, we conduct a detailed analysis of models’ potential in fact-checking real-world claims, including their accuracy and reliability.
3.1 Evaluation Prompt
We simultaneously obtain the predictions, explanations, and confidence levels from MLLMs with the prompt below. This prompt is formulated following the example of prompts found in related fact-checking work Pan et al. (2023); Min et al. (2023), while for acquiring explanations and confidence estimates we refer to Xiong et al. (2023).
Is it true that "CLAIM"? True or False? Use the following format to provide your answer:
Prediction: [True or False]
Explanation: [put your evidence and step-by-step reasoning here]
Confidence Level: [please show the percentage]
Note: The confidence level indicates the degree of certainty you have about your answer and is represented as a percentage. For instance, if your confidence level is , it means you are certain that your answer is correct and there is a chance that it may be incorrect.
3.2 Metrics
Response Types
Based on the three components (prediction, explanation, and confidence score) in the model’s response, we categorize them into four types: True, False, Uncertain, and Deny. Deny refers to instances where the model fails to respond according to our requirements, missing any component. Uncertain denotes cases where the model explicitly expresses uncertainty, with statements such as “cannot confirm”, “cannot verify”, or “cannot determine”, etc. Others are classified as True or False based on the model’s stated prediction. This classification scheme is designed to reveal the model’s inherent biases through the distribution of True and False responses, ability to follow instructions the number of Deny responses, and assess its knowledge-awareness through the frequency of Uncertain cases.
Accuracy Metrics
Assuming the total count of True (T), False (F), Uncertain (U) and Deny (D) responses is . Since not all responses include predictions, we denote the number of samples containing predictions as , with . We introduce two distinct accuracy metrics, True/False Accuracy is defined as , and Overall Accuracy is defined as , where represent the number of correctly predicted samples within the corresponding group. The first metric posits that predictions made with expressed uncertainty should be omitted from accuracy assessments. The second metric provides a holistic prediction-based score over the same set of data points for all models, thus allowing for a fair comparison.
3.3 Models
We use OpenAI’s API to collect responses from GPT-4V (gpt-4-1106-preview). We also experiment with open-source MLLMs, including LLaVA-1.5 (7b and 13b, Liu et al. 2023b, a), InstructBLIP (Flan-T5-xxl, Dai et al. 2023), and MiniGPT (v2, Chen et al. 2023). All parameters are set to the default values, with max_tokens=300 for GPT-4V. These models are notably representative and competitive in a multimodal setup.111We hereafter omit the version numbers of the open-source models, using LLaVA(7b), LLaVA(13b), MiniGPT, and InstructBLIP instead.
3.4 Datasets
Fauxtography |
Cosmos |
MOCHEG |
Post-4V |
True |
200 |
240 |
267 |
81 |
False |
200 |
240 |
237 |
105 |
Total |
400 |
480 |
504 |
186 |
Fauxtography
is a multimodal fact-checking dataset sourced from the websites Snopes and Reuters Zlatkova et al. (2019). Due to constraints posed by the GPT-4V API, such as decreased query speeds after exceeding a daily limit, we select a random subset of 400 entries, ensuring a balanced composition of 200 true and 200 false instances.
COSMOS
is a dataset built from Snopes and News Outlets, and tailored for training models to autonomously identify image and text pairs that are out of context Aneja et al. (2023), i.e. falsely paired. We randomly select 240 true and 240 false samples from the test split.
MOCHEG
is originally based on textual claims from Snopes and PolitiFact, with associated images serving as evidence Yao et al. (2023). This implies that a single claim can correspond to multiple images, many of which are not explicitly required for the claim to be verified. We first identify multimodal claims through keywords such as photograph, image, etc. Subsequently, we manually select images that were explicitly requested for verification and filter out those with overlaying words like fake, misleading, or miscaptioned. This procedure results in a total of 504 data entries.
Post-4V
To mitigate potential data leakage, we obtain new data from the Snopes based on articles published after the release of GPT-4V (Sept. 26, 2023). We adhere to the same filtering approach used for MOCHEG, resulting in 186 samples.
Details of all data samples used are presented in Table 1. Overall, the sample distribution is balanced, containing only two types of labels: True and False.222We exclude data labeled as Mixture, Unproven, and other labels that lack clear veracity. Additionally, we collected the URLs of the web pages corresponding to the claims and their fact-checking articles.
| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
Image |
Model |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
| Yes |
GPT |
158 |
195 |
29 |
18 |
81.9 |
82.1 |
179 |
204 |
83 |
14 |
86.2 |
80.0 |
216 |
223 |
37 |
28 |
87.2 |
83.5 |
54 |
98 |
26 |
8 |
79.6 |
73.7 |
LLa.(7b) |
337 |
6 |
1 |
56 |
53.4 |
46.3 |
449 |
7 |
4 |
20 |
52.0 |
50.0 |
409 |
6 |
0 |
89 |
54.2 |
44.8 |
157 |
1 |
3 |
25 |
42.0 |
37.1 |
LLa.(13b) |
286 |
98 |
3 |
13 |
54.4 |
52.3 |
331 |
134 |
13 |
2 |
69.7 |
67.7 |
363 |
112 |
28 |
1 |
61.5 |
58.1 |
116 |
65 |
2 |
3 |
56.3 |
55.9 |
MiniGPT |
55.5 |
62.1 |
54.3 |
55.9 |
Instruct. |
54.3 |
74.4 |
52.6 |
59.1 |
| No |
GPT |
105 |
202 |
32 |
61 |
80.1 |
71.0 |
111 |
253 |
50 |
66 |
64.0 |
57.5 |
153 |
265 |
38 |
48 |
73.9 |
68.9 |
35 |
68 |
10 |
73 |
74.8 |
67.7 |
LLa.(7b) |
398 |
0 |
2 |
0 |
50.3 |
50.0 |
474 |
0 |
4 |
2 |
50.2 |
49.8 |
502 |
0 |
2 |
0 |
53.0 |
52.8 |
181 |
0 |
4 |
1 |
43.1 |
43.0 |
LLa.(13b) |
167 |
214 |
17 |
2 |
62.5 |
62.0 |
263 |
195 |
20 |
2 |
76.6 |
75.8 |
168 |
316 |
20 |
0 |
62.0 |
61.5 |
40 |
124 |
22 |
0 |
60.4 |
60.8 |
MiniGPT |
56.8 |
60.4 |
54.0 |
58.1 |
Instruct. |
55.8 |
68.3 |
48.6 |
61.8 |
Yes |
Baseline |
50.0 |
50.0 |
53.0 |
56.5 |
4 Evaluation and Analysis
4.1 RQ1: Comprehensive Performance
Table 2 presents the comprehensive performance of various models across four datasets. The "Image" column clarifies the input type, with "Yes" indicating multimodal inputs and "No" signifying text-only inputs. MiniGPT and InstructBLIP, while incapable of providing explanations and confidence scores, still manage to generate predictions; thus, only the Overall Accuracy metric is assessed. The accuracy of the majority class is noted at the bottom of the table.
The upper section of Table 2 reveals that GPT-4V outperforms other models in both True/False and Overall Accuracy metrics, with True/False Accuracy frequently exceeding Overall Accuracy. It implies GPT-4V’s heightened precision when it responds with confidence. LLaVA’s accuracy consistently improves with increased model size. MiniGPT and InstructBLIP surpass the baseline and LLaVA(7b), nearing the accuracy levels of the 13b model. In addition, LLaVA(7b) shows a strong bias, almost always predicting True. As the model size increases, predictions become more balanced, and the expression of uncertainty rises. Furthermore, there is a significant reduction in the number of Deny responses, indicating that larger models are more capable of following instructions. The high Deny responses from GPT-4V are mainly due to its safety alignment, often replying with statements like "I’m sorry, but I cannot assist with this request." or "This content may violate our content policy."
Figure 2 displays multimodal claims and corresponding responses from GPT-4V and LLaVA(13b), with the annotations w. Image and w/o Image next to the model names indicating whether an image is in the input, respectively. From the perspective of explanations, GPT-4V’s responses contain a richer set of facts. For instance, it explicitly states, "there was never an official campaign for Michelle Obama in the 2020 election," in the 1st case. The 5th example provides detailed background information about Margaret Hamilton, including her role in "leading the development of onboard flight software for the Apollo mission." Moreover, GPT-4V adeptly incorporates other details from the images, such as referencing a "Google image search: Creepy Joe Biden" in the 4th image. In comparison, LLaVA(13b) lacks the capability to counteract manipulated images, as evidenced in the 2nd and 4th examples. Additionally, LLaVA’s explanations are typically brief, often limited to a description of the image followed by a conclusion about its veracity, indicating inefficient reasoning.
4.2 RQ2: Impact of Images: Multimodal VS. Text-Only inputs
By comparing data in Table 2, it is evident that images significantly impact the predictions of GPT-4V, with the model’s accuracy decreasing to various extents. Simultaneously, the instances where the model refuses to respond have noticeably increased. As in Figure 2, GPT-4V’s explanations indicate its awareness of missing images, using phrases like "there is no verified public record" or "without any provided context." This suggests that GPT-4V engages in knowledge retrieval to verify the existence or context of the images.
Counterintuitively, open-source models demonstrate higher accuracy with text-only inputs in many cases compared to multimodal inputs. To understand this, we scrutinize their explanations and find two specific scenarios. First, when the claim is false and the image is altered, as illustrated in the 1st and 4th examples in Figure 2, models can accurately predict the veracity of textual claims. However, when images are included, the models, failing to recognize the alterations, are misled into making incorrect predictions. Second, when both the image and the claim are true, as demonstrated in the 3rd example, models that consider images might deem the claim False due to inadequate evidence. In contrast, predictions for textual claims remain true, even with hallucinations highlighted in yellow. This implies that the models were not exposed to enough negative examples during the fine-tuning phase.
| Fauxtography | COSMOS | MOCHEG | Post-4V | |||||||||||||||||||||
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
EN |
158 |
195 |
29 |
18 |
81.9 |
82.1 |
179 |
204 |
83 |
14 |
86.2 |
80.0 |
216 |
223 |
37 |
28 |
87.2 |
83.5 |
54 |
98 |
26 |
8 |
79.6 |
73.7 |
ZH |
144 |
188 |
48 |
20 |
78.6 |
75.3 |
180 |
168 |
100 |
32 |
83.0 |
72.5 |
216 |
198 |
50 |
40 |
84.3 |
79.2 |
62 |
82 |
30 |
12 |
78.5 |
68.3 |
DE |
154 |
173 |
30 |
43 |
80.7 |
72.3 |
156 |
223 |
45 |
56 |
77.0 |
67.1 |
224 |
198 |
17 |
65 |
83.9 |
75.0 |
55 |
87 |
17 |
27 |
76.8 |
64.5 |
AR |
101 |
82 |
14 |
203 |
73.8 |
73.5 |
88 |
109 |
19 |
264 |
68.0 |
64.6 |
137 |
77 |
13 |
277 |
84.1 |
78.2 |
29 |
26 |
9 |
122 |
79.6 |
69.4 |
| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
idx |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
| GPT-4V |
0 |
158 |
195 |
29 |
18 |
81.9 |
82.1 |
179 |
204 |
83 |
14 |
86.2 |
80.0 |
216 |
223 |
37 |
28 |
87.2 |
83.5 |
54 |
98 |
26 |
8 |
79.6 |
73.7 |
2 |
158 |
181 |
32 |
29 |
84.1 |
81.3 |
163 |
219 |
75 |
23 |
85.1 |
80.8 |
205 |
234 |
34 |
31 |
82.1 |
81.2 |
55 |
95 |
22 |
14 |
77.8 |
75.3 |
5 |
143 |
202 |
31 |
24 |
81.2 |
76.3 |
175 |
216 |
72 |
17 |
88.0 |
81.5 |
179 |
250 |
44 |
31 |
78.6 |
75.6 |
51 |
106 |
18 |
11 |
76.4 |
71.5 |
| LLaVA |
0 |
286 |
98 |
3 |
13 |
54.4 |
52.3 |
331 |
134 |
13 |
2 |
69.7 |
67.7 |
363 |
112 |
28 |
1 |
61.5 |
58.1 |
116 |
65 |
2 |
3 |
56.3 |
55.9 |
1 |
197 |
163 |
7 |
33 |
48.6 |
48.5 |
169 |
236 |
15 |
60 |
69.6 |
66.3 |
247 |
211 |
9 |
37 |
57.6 |
57.9 |
68 |
105 |
5 |
8 |
56.6 |
57.5 |
2 |
170 |
223 |
4 |
3 |
49.1 |
49.5 |
197 |
274 |
7 |
2 |
69.0 |
68.5 |
318 |
178 |
7 |
1 |
58.1 |
57.9 |
41 |
141 |
4 |
0 |
56.0 |
57.0 |
3 |
243 |
145 |
8 |
4 |
52.6 |
53.0 |
243 |
222 |
13 |
2 |
71.8 |
70.2 |
295 |
197 |
11 |
1 |
57.3 |
56.3 |
82 |
93 |
10 |
1 |
56.5 |
56.5 |
4 |
211 |
180 |
6 |
3 |
62.1 |
62.0 |
232 |
244 |
3 |
1 |
77.3 |
77.1 |
210 |
286 |
4 |
4 |
57.3 |
56.3 |
55 |
129 |
2 |
0 |
57.6 |
57.0 |
5 |
241 |
154 |
3 |
2 |
63.3 |
63.0 |
313 |
163 |
4 |
0 |
80.5 |
80.4 |
260 |
236 |
8 |
0 |
62.1 |
62.3 |
109 |
75 |
2 |
0 |
52.7 |
52.7 |
4.3 RQ3: Robustness to Prompts
Multilingual prompts
Since our datasets primarily originate in English-language websites, multilingual prompts can assess the adaptability of intrinsic knowledge across different languages and mitigate data leakage. We translate our prompt template and claims into Chinese, German, and Arabic using the Google Translate API. Table 3 presents the performance across different languages. Notable variations can be observed in adherence to instructions, particularly in Arabic, which exhibits a higher frequency of Deny responses across datasets. Nevertheless, the overall accuracy levels remain comparable, suggesting that the responses’ lack of explanations and confidence scores in Arabic did not significantly impact model performance. Moreover, the True/False Accuracy generally surpasses the Overall Accuracy, indicating GPT-4V’s good knowledge awareness. We do not conduct experiments on open-source models as they are predominantly trained and tested in English.
Prompts with similar semantics
We employ ChatGPT to generate semantically similar yet distinct prompts, as depicted in Figure 3. The part pertaining to explanations and confidence scores is identical to the original prompt. We randomly select two prompts for testing GPT-4V and all generated prompts for the LLaVA(13b) model. The outcomes of the experiments are detailed in Table 4, where Index 0 denotes the original prompt, and the other indices represent the newly crafted prompts. The data reveals that GPT-4V maintains commendable consistency, while LLaVA demonstrates variable biases across different prompts. For the same dataset, such as Fauxtography, models tend to predict the majority of claims as True when using the original prompt (index 0). In contrast, the prompt of index 2 leads the models to predict most claims as False. Moreover, on Fauxtography and COSMOS, the maximum discrepancy in overall accuracy reached and , respectively, compared to just and for GPT-4V. It indicates that the LLaVA(13b) model is highly sensitive to prompts and fails to maintain consistency.
4.4 RQ4: Improving Capabilities via In-Context Learning
| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
Demo |
Num |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
T |
F |
U |
D |
T/F |
O |
0 |
0 |
286 |
98 |
16 |
0 |
54.4 |
52.8 |
331 |
134 |
14 |
1 |
69.7 |
67.7 |
363 |
112 |
29 |
0 |
61.5 |
58.1 |
116 |
65 |
5 |
0 |
56.4 |
55.9 |
| 1 |
1 |
244 |
135 |
9 |
12 |
60.9 |
61.8 |
295 |
171 |
7 |
7 |
75.3 |
74.2 |
255 |
229 |
12 |
8 |
62.6 |
62.7 |
95 |
88 |
3 |
0 |
56.3 |
55.4 |
2 |
182 |
161 |
6 |
51 |
59.5 |
60.0 |
234 |
220 |
8 |
18 |
75.8 |
73.5 |
169 |
282 |
10 |
43 |
57.4 |
55.5 |
42 |
109 |
6 |
29 |
67.5 |
63.4 |
| 2 |
1 |
264 |
121 |
4 |
11 |
61.6 |
62.8 |
337 |
131 |
7 |
5 |
73.9 |
73.5 |
314 |
180 |
4 |
6 |
64.8 |
65.1 |
97 |
79 |
10 |
0 |
57.4 |
55.9 |
2 |
176 |
204 |
10 |
10 |
60.0 |
60.5 |
233 |
230 |
11 |
6 |
80.3 |
79.6 |
199 |
291 |
8 |
6 |
61.2 |
60.7 |
44 |
131 |
11 |
0 |
60.0 |
60.8 |
| 3 |
1 |
211 |
185 |
1 |
3 |
62.6 |
62.8 |
279 |
195 |
6 |
0 |
72.2 |
71.7 |
236 |
266 |
2 |
0 |
61.8 |
61.5 |
75 |
111 |
0 |
0 |
60.2 |
60.2 |
2 |
191 |
198 |
9 |
2 |
64.1 |
63.8 |
263 |
201 |
16 |
0 |
75.4 |
75.4 |
206 |
288 |
9 |
1 |
61.1 |
61.7 |
54 |
127 |
4 |
1 |
59.7 |
59.7 |
| 4 |
1 |
191 |
193 |
9 |
7 |
62.2 |
62.0 |
260 |
203 |
16 |
1 |
78.0 |
77.9 |
210 |
283 |
11 |
0 |
60.6 |
60.9 |
50 |
126 |
8 |
2 |
60.8 |
60.8 |
2 |
196 |
188 |
7 |
9 |
62.8 |
62.8 |
259 |
207 |
12 |
2 |
77.5 |
77.5 |
207 |
284 |
11 |
2 |
61.1 |
61.3 |
52 |
122 |
10 |
2 |
61.5 |
61.8 |
As demonstrated in the 1st and 4th examples in Figure 2, the LLaVA(13b) model struggles to distinguish claims associated with manipulated images. The image content is consistent with the textual claim but contradicts real-world facts. To enhance the model’s comprehension of the task, checking facts rather than the consistency between the image and text, we experiment with employing in-context learning to improve its performance. In detail, we obtain GPT-4V’s responses for claims outside our test datasets and employ these as instructive examples for LLaVA(13b). Due to LLaVA’s limitation with multiple image inputs, we combine different images vertically with a 20-pixel gap between them. The linkage between images and claims is narratively indicated, as illustrated in the following prompt:
Use the following format to answer whether the claim regarding the pictures is True or False:
SAME FORMAT AS IN THE ORIGINAL PROMPT.
For the first picture, claim: "CLAIM1"
Prediction: False
Explanation: XXX
Confidence Level: XXX
For the second picture, claim: "CLAIM2"
Prediction: True
Explanation: XXX
Confidence Level: XXX
For the third picture, claim: "CLAIM"
We construct four sets of demonstrations, each comprising one true and one false claim. We report the performances when using only the false claim and employing both claims as examples. We observe a universal improvement in model accuracy when examples are presented in Table 5. However, the use of the second example does not contribute to accuracy. In the explanations provided by LLaVA, we can observe two effects: one is the average length of the explanation. For instance, on MOCHEG, with the presentation of 0, 1, and 2 examples from the first set, the average sentence length of the model’s explanations are 3.09, 4.02, and 3.75, respectively. Another change is the frequency of the term "manipulated," which increases from 4 to 170 and 193 on MOCHEG. This suggests that the model can learn to analyze from the perspective of image manipulation. The last two rows of Figure 2 display responses from the LLaVA(13b) model. In the 4th example, it is evident that through learning, the model recognizes the altered part in the image and explains the motive behind the modifications: "but the gun and menacing pose are added elements," and "image is likely created to spread misinformation and discredit Joe Biden or to provoke a reaction from viewers." However, some samples are not correctly identified, such as the 1st example, and we observe that LLaVA produces outdated information, as highlighted in green.
4.5 RQ5: Calibration Evaluation
Calibration refers to the alignment between a model’s predicted probabilities (or confidence levels) and the actual outcomes. It reflects the model’s knowledge-awareness and ability to accurately estimate the confidence in its responses Geng et al. (2023). In this evaluation, we use verbalized confidence, which is readily accessible. Figure 4 displays the confidence distributions alongside calibration curves for GPT-4V and LLaVA(13b), breaking down the data into 10% confidence intervals. For instance, a value of 95 corresponds to confidence levels within the bracket. GPT-4V’s confidence scores are largely clustered in the interval, whereas LLaVA(13b)’s scores are more frequently found in the range. The calibration curves reveal that both models exhibit a positive correlation between confidence levels and accuracy, with GPT-4V’s calibration closely mirroring the ideal curve (illustrated by the dashed line), indicating well-calibrated confidence scores. Conversely, LLaVA(13b)’s curve suggests a propensity for overconfidence. The Expected Calibration Errors (ECEs, Guo et al. 2017), detailed in Table 6, further substantiate that GPT-4V maintains superior calibration.
ECE |
Fauxtography |
COSMOS |
MOCHEG |
Post-4V |
GPT-4V |
7.76 |
4.36 |
4.88 |
11.83 |
LLaVA |
30.20 |
15.26 |
22.82 |
28.45 |
4.6 RQ6: Taxonomy of Failure Reasons
Upon examining GPT-4V’s explanations, we identify six primary categories of failure reasons. Failures attributed to Lack of Knowledge occur when responses are based on trivial or incomplete information, often contradicting common sense. Recognition Errors arise from the model’s visual module misinterpreting images, including failing to identify altered or AI-generated visuals. Reasoning Errors occur when the model incorrectly interprets images due to factors like unnatural distortions or unusual perspectives, or there are self-contradictions in reasoning. Failures due to Partial Facts happen when the model addresses only a subset of the facts present in a textual claim. Image Issues involve failures related to poor image quality, such as blurriness or incompleteness, affecting the model’s ability to recognize content accurately. Ambiguity Claims involve semantic uncertainties, where a term could reference multiple entities. Examples beyond the lack of knowledge are presented in Figure 5. In subfigure (a), GPT-4V misinterprets scars from a dog’s burn as being painted or digitally altered. In subfigure (b), it correctly identifies the locations of Notre Dame and the Eiffel Tower but engages in erroneous reasoning. In subfigure (c), the model understands that a kiss can release hormones like oxytocin and knows that MR images cannot capture chemical reactions, with its predictions primarily influenced by the first fact. In subfigure (d), the model cannot discern that the scene depicted is related to posture training based on the image’s content. In subfigure (e), "Playboy" has multiple meanings, which could refer to a publisher or magazine’s name. We also manually tall the frequency of these errors across various datasets, as detailed in Table 7. Overall, the primary cause of errors is attributed to Lack of Knowledge, suggesting that augmenting the system with additional evidence could enhance its performance. There is room for improving the model’s ability to understand and reason about AI-generated or distorted images.
Know |
Recog |
Reas |
Partial |
Image |
Ambi |
| Fauxtography |
59.3 |
20.3 |
17.2 |
1.6 |
1.6 |
0.0 |
| COSMOS |
80.3 |
5.9 |
9.8 |
2.0 |
0 |
2.0 |
| MOCHEG |
73.2 |
3.6 |
7.1 |
3.6 |
5.4 |
7.1 |
| Post-4V |
85.2 |
11.1 |
0.0 |
0.0 |
0.0 |
3.7 |
We also explore the utilization of GPT-4 for detecting nonfactual content within explanations when fact-checking articles are available. We design a prompt comprising two steps: removing sentences containing no facts and predicting whether the reference article supports the sentence. Due to the limited content space of the article, we do not present the specific prompt. Figure 6 showcases two instances where, although the predictions are correct, the facts within the explanations are inaccurate. This underscores the necessity for further examination of their veracity. The photograph on the left, featuring a young girl named Fouzia, was captured in Morocco by Giacomo Pirozzi, a photographer for the United Nations Children’s Fund (UNICEF). However, GPT-4V incorrectly attributes it to photographer Terri Gold in Yemen, 1998. The right photo is incorrectly identified as from the movie "Platoon," actually, it is from the movie "Tropic Thunder." The returned sentences also contained substantial image descriptions not present in the original fact-checking articles. Improving the automatic assessment of the factuality in explanations is a focus for our future work.
5 Conclusion and Future Work
We investigated the capabilities of MLLMs to fact-check real-world claims, relying solely on their parametric knowledge and reasoning capabilities without external references. We proposed an evaluation framework, designed various experiments to address the research questions. Our results indicated that state-of-the-art MLLMs, such as GPT-4V, have the potential to assist professional fact-checkers: they can enhance their efficiency by providing reference predictions, valuable clues, and explanatory insights, together with confidence. Open-source models exhibit sensitivity to text, with different prompts revealing various biases. Moreover, they are susceptible to being deceived by altered images, and even in the absence of images, they tend to produce hallucinations pertaining to image descriptions. Our work demonstrates both the potential and limitations of MLLMs, laying the groundwork for further leveraging MLLMs to support fact-checking.
In future work, we plan to investigate how to enhance the fact-checking capabilities of models, focusing not only on improving accuracy but also on strengthening robustness and augmenting the knowledge reasoning and the verification abilities required for fact-checking. We further plan to explore how MLLMs can better support fact-checking when connected to external knowledge.

### v2

@@
 Abstract
-Multimodal large language models (MLLMs) carry the potential to support humans in processing vast amounts of information.
-While MLLMs are already being used as a fact-checking tool, their abilities and limitations in this regard are understudied.
+Misinformation poses a real-world threat, especially when combined with misleading images.
+Multimodal large language models (MLLMs), which combine image comprehension with the rich knowledge and explanatory capability of language models, have become tools for humans to process vast information.
+However, their capacity and limitations as multimodal tools in assisting with fact-checking remain understudied.
 Here is aim to bridge this gap.
@@
 Our methodology is evidence-free, leveraging only these models’ intrinsic knowledge and reasoning capabilities.
-By designing prompts that extract models’ predictions, explanations, and confidence levels, we delve into research questions concerning model accuracy, robustness, and reasons for failure.
-We empirically find that (1) GPT-4V exhibits superior performance in identifying malicious and misleading multimodal claims, with the ability to explain the unreasonable aspects and underlying motives, and (2) existing open-source models exhibit strong biases and are highly sensitive to the prompt.
-Our study offers insights into combating false multimodal information and building secure, trustworthy multimodal models.
-To the best of our knowledge, we are the first to evaluate MLLMs for real-world fact-checking.
+By designing prompts that extract models’ predictions, explanations, and confidence levels, we conduct a detailed analysis of the model’s accuracy, bias, and other key factors.
+We empirically find that (1) GPT-4V exhibits surprising performance across various datasets, with an accuracy rate exceeding , and is capable of providing impressive explanations, and (2) even with the aid of prompt ensembles and in-context learning, open-source models significantly lag in performance.
+However, they show potential in remembering checked claims and reasoning out manipulated images.
+We also summarize the failure reasons, which contributes to strategies for future improvements.
+Our study offers insights into leveraging MLLMs to combat multimodal misinformation.
 1 Introduction
-Misinformation is a significant challenge on the internet, especially with regard to multimodal claims, which combine text, images, videos, and other media types Mubashara et al.
+Misinformation is a significant challenge on the internet, especially with regard to multimodal claims, which combine text, images, videos, and other media types Akhtar et al.
 (2023).
-The visual component in these cases could be manipulated, used out-of-context, or accompanied by text that makes a false claim Huh et al.
+The visual component in these cases could be manipulated or used out-of-context (OOC) to make a false claim Huh et al.
 (2018); Luo et al.
@@
 In such cases, fact-checkers and the tools they employ need to be able to handle multiple modalities.
-Large language models (LLMs) store extensive information about the world, which exceeds what any individual human can know and is accessible in a far more human-friendly way than any search engine can offer Zhao et al.
+Large language models (LLMs) store extensive information, which exceeds what any individual human can know and is accessible in a far more human-friendly way than any search engine can offer Zhao et al.
 (2023); Hu et al.
-(2023a).
+(2024).
 As such, they can be a powerful tool in the hands of fact-checkers, who often seek extra factual knowledge to verify a claim Nakov et al.
@@
 Yet, this pipeline approach requires an extra step of evidence retrieval, which is computationally heavy and error-prone, while it overlooks the vast amount of knowledge stored in the parameters of LLMs Hu et al.
-(2023b).
-Additionally, LLMs are now giving way from text-only to multimodal models, which can process signals from multiple sources and perform a wider range of tasks Liu et al.
-(2023b); Zhang et al.
-(2023); Yin et al.
 (2023).
-We hypothesize that multimodal large language models (MLLMs) trained on large amounts of data can serve as a sufficient substitute for retrieving evidence and perform fact-checking in an evidence-free fashion, i.e., relying solely on their parametric knowledge.
-Despite their great promise, however, MLLMs have limitations as well, specifically with respect to factuality Fu et al.
+We hypothesize that multimodal large language models (MLLMs) trained on large amounts of data can serve as a sufficient substitute for the retrieval of evidence and perform fact-checking in an evidence-free fashion, i.e., relying solely on their parametric knowledge.
+Despite their great promise, MLLMs have limitations as well, specifically with respect to factuality Fu et al.
 (2023); Liu et al.
-(2023c); Xu et al.
-(2023).
+(2023c).
 Models may produce hallucinations Huang et al.
@@
 (2023b) and suffer from a lack of knowledge or exhibit biases Hu et al.
-(2023a); Wang et al.
+(2024); Wang et al.
 (2023).
@@
 Based on the data obtained, we can further assess the models’ capabilities from various perspectives, addressing the following research questions:
-RQ1: How accurate are MLLMs in identifying misinformation in multimodal claims?
-RQ2: How do different modalities (text and image) contribute to the assessment of multimodal content?
-RQ3: Are models sensitive to prompts in different languages or expressions, or do they exhibit robustness?
-RQ4: Can in-context learning enhance model performance?
-RQ5: Can MLLMs accurately estimate the correctness of their responses in terms of a confidence score?
-RQ6: What types of errors are most common for MLLMs employed as multimodal fact-checkers?
-Our investigation spans an array of vision-language models, including GPT-4V, LLaVA-1.5 Liu et al.
-(2023b), MiniGPT-v2 Chen et al.
-(2023), and InstructBLIP Dai et al.
-(2023), which stand as leaders in this domain.
-We evaluate these models across three existing multimodal fact-checking datasets Zlatkova et al.
-(2019); Aneja et al.
-(2023); Yao et al.
-(2023), and one new dataset, which we build from recent check-worthy claims to mitigate the risk of data contamination.
+RQ1 Performance Evaluation: How good are MLLMs in identifying multimodal misinformation? Can they perform accurate reasoning?
+RQ2 Improving Approaches: How can we effectively boost the model’s fact-checking capabilities? Can the improvements be reflected in the reasoning, beyond accuracy metrics?
+RQ3 Taxonomy of Failure Reasons: What are the typical errors for MLLMs employed as multimodal fact-checkers?
+Our investigation spans an array of vision-language models, including GPT-4V, MiniGPT-v2 Chen et al.
+(2023) and LLaVA-1.5 Liu et al.
+(2023b), which stand as leaders in this domain.
+We evaluate these models across three existing multimodal fact-checking datasets: Fauxtography Zlatkova et al.
+(2019), COSMOS Aneja et al.
+(2023), MOCHEG Yao et al.
+(2023), and one new dataset, which build from recent check-worthy claims to mitigate the risk of data contamination.
+We use prompt ensembles (PE) and in-context learning (ICL) to help improve the open-source MLLMs.
+We do not use fine-tuning because models are required to explain its reasoning process and provide confidence scores, while high-quality annotated data is scarce.
 Our study finds that MLLMs display a range of fact-checking capabilities.
-GPT-4V, in particular, excels in our tests, exhibiting high accuracy, useful explanations, and good calibration, with the accuracy on different datasets generally reaching around .
-On the other hand, open-source models like MiniGPT-v2 and InstructBLIP lack explanatory and uncertainty-reporting capabilities.
-LLaVA models show a strong sensitivity to prompt formulation, a tendency to comply with the question in the prompt, and hallucinatory behaviors.
+GPT-4V, in particular, excels in our tests, exhibiting high accuracy, useful explanations, and good calibration.
+The accuracy on various datasets generally reaches around .
+MiniGPT-v2 lacks explanatory and uncertainty-reporting capabilities.
+LLaVAs can follow instructions, and both PE and ICL enhance their performance to varying degrees, with the latter providing greater improvement.
+On Fauxtography and COSMOS, ICL increases the Overall Accuracy from 52.3% to 62.3% and from 67.7% to 76.5%, respectively.
+The majority baseline for both datasets is 50%.
 In summary, we systematically evaluate existing multimodal models to determine their capability to support real-world fact-checking.
-We formulate several basic research questions and assess their accuracy, robustness, calibration, and reasons for failure across various datasets.
-Our work fills a research gap, providing a comprehensive overview of the domain and shedding light on both fact-checking and MLLMs’ factuality.
-As far as we are aware, our study is the first assessment of MLLMs in the context of real-world fact-checking.
-2 Background and Related Work
+Our approach is more comprehensive than prior studies, which are limited to specific aspects such as face spoofing, forgery detection, or out-of-context detection Shi et al.
+(2024); Qi et al.
+(2024).
+We formulate several research questions to evaluate these models based on accuracy, bias, reasoning capabilities, and errors across a variety of datasets.
+Our work fills a research gap, providing a comprehensive overview of the domain and shedding light on leveraging MLLMs to combat multimodal misinformation.
+2 Related Work
 LLMs for Text-Only Fact-Checking
-Hoes et al.
-(2023) find that ChatGPT accurately classifies 69% of statements in a dataset built from the fact-checking platform PolitiFact.
-FactLLaMA Cheung and Lam (2023) integrates external evidence into the instruct-tuning process to enhance the model’s ability to leverage evidence for predictions.
+Caramancion (2023) compared the performance of models such as ChatGPT 3.5 and ChatGPT 4.0 in news fact-checking.FactLLaMA Cheung and Lam (2023) integrates external evidence into the instruct-tuning process to enhance the model’s ability to leverage evidence for predictions.
 Studies have also focused on leveraging the capabilities of LLMs to develop automated frameworks for decomposing claims, generating queries, and predicting based on gathered evidence, thereby augmenting the interpretability of the fact-checking process Pan et al.
@@
 Multimodal Fact-Checking
-Multimodal misinformation has been found to spread faster than its text-only counterparts Li and Xie (2020).
-It primarily falls into two categories: one involves textual claims about manipulated content across different modalities, while the other pertains to out-of-context misinformation, featuring unaltered images, audio, or videos in misleading contexts Luo et al.
+Multimodal misinformation primarily falls into two categories: one involves textual claims about manipulated content across different modalities, while the other pertains to out-of-context misinformation, featuring unaltered images, audio, or videos in misleading contexts Luo et al.
 (2021); Aneja et al.
@@
 (2023).
-The multimodal fact-checking is gnerally structured into distinct stages within pipelines, including claim detection extraction, evidence retrieval, verdict prediction, and explanation Nakov et al.
-(2021); Mubashara et al.
-(2023).
-In our work, we posit that evidence is parametrically encoded within the models, enabling us to bypass the evidence retrieval phase and directly engage in predictions and explanations for the targeted claims.
-Evaluation of MLLMs
-With the advancement of MLLMs, the systematic evaluation of their capabilities has become increasingly critical.
-A broad array of benchmarks has been established to estimate the MLLM’s performance across a variety of tasks, including perceptual tasks such as image recognition, image description, and optical character recognition, as well as cognitive tasks like commonsense reasoning and numerical calculation Yin et al.
-(2023); Fu et al.
-(2023); Liu et al.
-(2023c); Xu et al.
-(2023); Tong et al.
-(2024).
-Tu et al.
-(2023) access the resilience of MLLMs against various image-related attacks from a security perspective.
-However, there is an absence of research on evaluating the ability to conduct multimodal fact-checking with MLLMs.
-Our work aims to bridge this gap.
+Shi et al.
+(2024) proposed a benchmark to evaluate the performance of MLLMs in detecting face spoofing and forgery.
+Qi et al.
+(2024) introduced Sniffer, an MLLM designed for OOC detection and explanation, utilizing fine-tuning to boost the accuracy in identifying celebrities and clarifying inconsistencies between text and images.
+In our work, we posit that evidence is parametrically encoded within the models, enabling us to bypass the evidence retrieval phase and directly engage in predictions and explanations for the real-world claims.
 3 Evaluation Framework
-We propose an evaluation framework that analyzes model capabilities along the various dimensions illustrated in Figure 1.
-Our experiments are conducted in an evidence-free setting, meaning that models have access only to the multimodal claims without any external knowledge.
-Through extensive controlled experiments, we conduct a detailed analysis of models’ potential in fact-checking real-world claims, including their accuracy and reliability.
-3.1 Evaluation Prompt
-We simultaneously obtain the predictions, explanations, and confidence levels from MLLMs with the prompt below.
-This prompt is formulated following the example of prompts found in related fact-checking work Pan et al.
-(2023); Min et al.
-(2023), while for acquiring explanations and confidence estimates we refer to Xiong et al.
-(2023).
-Is it true that "CLAIM"? True or False? Use the following format to provide your answer:
-Prediction: [True or False]
-Explanation: [put your evidence and step-by-step reasoning here]
-Confidence Level: [please show the percentage]
-Note: The confidence level indicates the degree of certainty you have about your answer and is represented as a percentage.
-For instance, if your confidence level is , it means you are certain that your answer is correct and there is a chance that it may be incorrect.
-3.2 Metrics
-Response Types
-Based on the three components (prediction, explanation, and confidence score) in the model’s response, we categorize them into four types: True, False, Uncertain, and Deny.
-Deny refers to instances where the model fails to respond according to our requirements, missing any component.
-Uncertain denotes cases where the model explicitly expresses uncertainty, with statements such as “cannot confirm”, “cannot verify”, or “cannot determine”, etc.
-Others are classified as True or False based on the model’s stated prediction.
-This classification scheme is designed to reveal the model’s inherent biases through the distribution of True and False responses, ability to follow instructions the number of Deny responses, and assess its knowledge-awareness through the frequency of Uncertain cases.
-Accuracy Metrics
-Assuming the total count of True (T), False (F), Uncertain (U) and Deny (D) responses is .
-Since not all responses include predictions, we denote the number of samples containing predictions as , with .
-We introduce two distinct accuracy metrics, True/False Accuracy is defined as , and Overall Accuracy is defined as , where represent the number of correctly predicted samples within the corresponding group.
-The first metric posits that predictions made with expressed uncertainty should be omitted from accuracy assessments.
-The second metric provides a holistic prediction-based score over the same set of data points for all models, thus allowing for a fair comparison.
-3.3 Models
-We use OpenAI’s API to collect responses from GPT-4V (gpt-4-1106-preview).
-We also experiment with open-source MLLMs, including LLaVA-1.5 (7b and 13b, Liu et al.
-2023b, a), InstructBLIP (Flan-T5-xxl, Dai et al.
-2023), and MiniGPT (v2, Chen et al.
-2023).
-All parameters are set to the default values, with max_tokens=300 for GPT-4V.
-These models are notably representative and competitive in a multimodal setup.111We hereafter omit the version numbers of the open-source models, using LLaVA(7b), LLaVA(13b), MiniGPT, and InstructBLIP instead.
+We propose an evaluation framework that includes datasets, prompts, and evaluation metrics to address the research questions.
 3.1 Datasets
-Fauxtography |
-Cosmos |
-MOCHEG |
-Post-4V |
-True |
-200 |
-240 |
-267 |
-81 |
-False |
-200 |
-240 |
-237 |
-105 |
-Total |
-400 |
-480 |
-504 |
-186 |
+Due to constraints posed by the GPT-4V API, such as decreased query speeds after exceeding a daily limit, we sample data from various datasets and preprocess them to fit our evaluation framework.
+Additionally, we develop a new dataset to ensure the model has not encountered these data.
 Fauxtography
@@
 (2019).
-Due to constraints posed by the GPT-4V API, such as decreased query speeds after exceeding a daily limit, we select a random subset of 400 entries, ensuring a balanced composition of 200 true and 200 false instances.
+We select a random subset of 400 entries, ensuring a balanced composition of 200 true and 200 false instances.
 COSMOS
-is a dataset built from Snopes and News Outlets, and tailored for training models to autonomously identify image and text pairs that are out of context Aneja et al.
-(2023), i.e.
-falsely paired.
+is a dataset built from Snopes and News Outlets, and tailored for OOC detection Aneja et al.
+(2023).
 We randomly select 240 true and 240 false samples from the test split.
+Each image in this dataset comes with two captions, and we manually select one caption per image to ensure it aligns with our criteria.
 MOCHEG
@@
 Post-4V
-To mitigate potential data leakage, we obtain new data from the Snopes based on articles published after the release of GPT-4V (Sept.
-26, 2023).
-We adhere to the same filtering approach used for MOCHEG, resulting in 186 samples.
-Details of all data samples used are presented in Table 1.
-Overall, the sample distribution is balanced, containing only two types of labels: True and False.222We exclude data labeled as Mixture, Unproven, and other labels that lack clear veracity.
-Additionally, we collected the URLs of the web pages corresponding to the claims and their fact-checking articles.
-| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
-Image |
-Model |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-| Yes |
-GPT |
-158 |
-195 |
-29 |
-18 |
-81.9 |
-82.1 |
-179 |
-204 |
-83 |
-14 |
-86.2 |
-80.0 |
-216 |
-223 |
-37 |
-28 |
-87.2 |
-83.5 |
-54 |
-98 |
-26 |
-8 |
-79.6 |
-73.7 |
-LLa.(7b) |
-337 |
-6 |
-1 |
-56 |
-53.4 |
-46.3 |
-449 |
-7 |
-4 |
-20 |
-52.0 |
-50.0 |
-409 |
-6 |
-0 |
-89 |
-54.2 |
-44.8 |
-157 |
-1 |
-3 |
-25 |
-42.0 |
-37.1 |
-LLa.(13b) |
-286 |
-98 |
-3 |
-13 |
-54.4 |
-52.3 |
-331 |
-134 |
-13 |
-2 |
-69.7 |
-67.7 |
-363 |
-112 |
-28 |
-1 |
-61.5 |
-58.1 |
-116 |
-65 |
-2 |
-3 |
-56.3 |
-55.9 |
-MiniGPT |
-55.5 |
-62.1 |
-54.3 |
-55.9 |
-Instruct. |
-54.3 |
-74.4 |
-52.6 |
-59.1 |
-| No |
-GPT |
-105 |
-202 |
-32 |
-61 |
-80.1 |
-71.0 |
-111 |
-253 |
-50 |
-66 |
-64.0 |
-57.5 |
-153 |
-265 |
-38 |
-48 |
-73.9 |
-68.9 |
-35 |
-68 |
-10 |
-73 |
-74.8 |
-67.7 |
-LLa.(7b) |
-398 |
-0 |
-2 |
-0 |
-50.3 |
-50.0 |
-474 |
-0 |
-4 |
-2 |
-50.2 |
-49.8 |
-502 |
-0 |
-2 |
-0 |
-53.0 |
-52.8 |
-181 |
-0 |
-4 |
-1 |
-43.1 |
-43.0 |
-LLa.(13b) |
-167 |
-214 |
-17 |
-2 |
-62.5 |
-62.0 |
-263 |
-195 |
-20 |
-2 |
-76.6 |
-75.8 |
-168 |
-316 |
-20 |
-0 |
-62.0 |
-61.5 |
-40 |
-124 |
-22 |
-0 |
-60.4 |
-60.8 |
-MiniGPT |
-56.8 |
-60.4 |
-54.0 |
-58.1 |
-Instruct. |
-55.8 |
-68.3 |
-48.6 |
-61.8 |
-Yes |
-Baseline |
-50.0 |
-50.0 |
-53.0 |
-56.5 |
-4 Evaluation and Analysis
-4.1 RQ1: Comprehensive Performance
-Table 2 presents the comprehensive performance of various models across four datasets.
-The "Image" column clarifies the input type, with "Yes" indicating multimodal inputs and "No" signifying text-only inputs.
-MiniGPT and InstructBLIP, while incapable of providing explanations and confidence scores, still manage to generate predictions; thus, only the Overall Accuracy metric is assessed.
-The accuracy of the majority class is noted at the bottom of the table.
-The upper section of Table 2 reveals that GPT-4V outperforms other models in both True/False and Overall Accuracy metrics, with True/False Accuracy frequently exceeding Overall Accuracy.
-It implies GPT-4V’s heightened precision when it responds with confidence.
-LLaVA’s accuracy consistently improves with increased model size.
-MiniGPT and InstructBLIP surpass the baseline and LLaVA(7b), nearing the accuracy levels of the 13b model.
-In addition, LLaVA(7b) shows a strong bias, almost always predicting True.
-As the model size increases, predictions become more balanced, and the expression of uncertainty rises.
-Furthermore, there is a significant reduction in the number of Deny responses, indicating that larger models are more capable of following instructions.
-The high Deny responses from GPT-4V are mainly due to its safety alignment, often replying with statements like "I’m sorry, but I cannot assist with this request." or "This content may violate our content policy."
-Figure 2 displays multimodal claims and corresponding responses from GPT-4V and LLaVA(13b), with the annotations w.
-Image and w/o Image next to the model names indicating whether an image is in the input, respectively.
-From the perspective of explanations, GPT-4V’s responses contain a richer set of facts.
-For instance, it explicitly states, "there was never an official campaign for Michelle Obama in the 2020 election," in the 1st case.
-The 5th example provides detailed background information about Margaret Hamilton, including her role in "leading the development of onboard flight software for the Apollo mission." Moreover, GPT-4V adeptly incorporates other details from the images, such as referencing a "Google image search: Creepy Joe Biden" in the 4th image.
-In comparison, LLaVA(13b) lacks the capability to counteract manipulated images, as evidenced in the 2nd and 4th examples.
-Additionally, LLaVA’s explanations are typically brief, often limited to a description of the image followed by a conclusion about its veracity, indicating inefficient reasoning.
-4.2 RQ2: Impact of Images: Multimodal VS. Text-Only inputs
-By comparing data in Table 2, it is evident that images significantly impact the predictions of GPT-4V, with the model’s accuracy decreasing to various extents.
-Simultaneously, the instances where the model refuses to respond have noticeably increased.
-As in Figure 2, GPT-4V’s explanations indicate its awareness of missing images, using phrases like "there is no verified public record" or "without any provided context." This suggests that GPT-4V engages in knowledge retrieval to verify the existence or context of the images.
-Counterintuitively, open-source models demonstrate higher accuracy with text-only inputs in many cases compared to multimodal inputs.
-To understand this, we scrutinize their explanations and find two specific scenarios.
-First, when the claim is false and the image is altered, as illustrated in the 1st and 4th examples in Figure 2, models can accurately predict the veracity of textual claims.
-However, when images are included, the models, failing to recognize the alterations, are misled into making incorrect predictions.
-Second, when both the image and the claim are true, as demonstrated in the 3rd example, models that consider images might deem the claim False due to inadequate evidence.
-In contrast, predictions for textual claims remain true, even with hallucinations highlighted in yellow.
-This implies that the models were not exposed to enough negative examples during the fine-tuning phase.
+We collect new data from Snopes, focusing on articles published after the release of GPT-4V (September 26, 2023).
+We apply the same filtering methodology used in MOCHEG, resulting in a dataset of 186 samples.
+| Fauxtography | COSMOS | MOCHEG | Post-4V | |
+| True | 200 | 240 | 267 | 81 |
+| False | 200 | 240 | 237 | 105 |
+| Total | 400 | 480 | 504 | 186 |
+3.2 Evaluation Prompt
+We simultaneously obtain the predictions, explanations, and confidence levels from MLLMs with the prompt below.
+This prompt is formulated following the example of prompts found in related fact-checking work Pan et al.
+(2023); Min et al.
+(2023), while for acquiring explanations and confidence estimates we refer to Xiong et al.
+(2024).
+We collect their verbalized confidence as it is increasingly used for decision-making during human-machine collaboration Geng et al.
+(2023).
+{adjustwidth}0.5cm0.5cm
+Is it true that "CLAIM"? True or False? Use the following format to provide your answer:
+Prediction: [True or False]
+Explanation: [put your evidence and step-by-step reasoning here]
+Confidence Level: [please show the percentage]
+Note: The confidence level indicates the degree of certainty you have about your answer and is represented as a percentage.
+For instance, if your confidence level is , it means you are certain that your answer is correct and there is a chance that it may be incorrect.
 | Fauxtography | COSMOS | MOCHEG | Post-4V | |||||||||||||||||||||
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-EN |
-158 |
-195 |
-29 |
-18 |
-81.9 |
-82.1 |
-179 |
-204 |
-83 |
-14 |
-86.2 |
-80.0 |
-216 |
-223 |
-37 |
-28 |
-87.2 |
-83.5 |
-54 |
-98 |
-26 |
-8 |
-79.6 |
-73.7 |
-ZH |
-144 |
-188 |
-48 |
-20 |
-78.6 |
-75.3 |
-180 |
-168 |
-100 |
-32 |
-83.0 |
-72.5 |
-216 |
-198 |
-50 |
-40 |
-84.3 |
-79.2 |
-62 |
-82 |
-30 |
-12 |
-78.5 |
-68.3 |
-DE |
-154 |
-173 |
-30 |
-43 |
-80.7 |
-72.3 |
-156 |
-223 |
-45 |
-56 |
-77.0 |
-67.1 |
-224 |
-198 |
-17 |
-65 |
-83.9 |
-75.0 |
-55 |
-87 |
-17 |
-27 |
-76.8 |
-64.5 |
-AR |
-101 |
-82 |
-14 |
-203 |
-73.8 |
-73.5 |
-88 |
-109 |
-19 |
-264 |
-68.0 |
-64.6 |
-137 |
-77 |
-13 |
-277 |
-84.1 |
-78.2 |
-29 |
-26 |
-9 |
-122 |
-79.6 |
-69.4 |
-| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
-idx |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-| GPT-4V |
-0 |
-158 |
-195 |
-29 |
-18 |
-81.9 |
-82.1 |
-179 |
-204 |
-83 |
-14 |
-86.2 |
-80.0 |
-216 |
-223 |
-37 |
-28 |
-87.2 |
-83.5 |
-54 |
-98 |
-26 |
-8 |
-79.6 |
-73.7 |
-2 |
-158 |
-181 |
-32 |
-29 |
-84.1 |
-81.3 |
-163 |
-219 |
-75 |
-23 |
-85.1 |
-80.8 |
-205 |
-234 |
-34 |
-31 |
-82.1 |
-81.2 |
-55 |
-95 |
-22 |
-14 |
-77.8 |
-75.3 |
-5 |
-143 |
-202 |
-31 |
-24 |
-81.2 |
-76.3 |
-175 |
-216 |
-72 |
-17 |
-88.0 |
-81.5 |
-179 |
-250 |
-44 |
-31 |
-78.6 |
-75.6 |
-51 |
-106 |
-18 |
-11 |
-76.4 |
-71.5 |
-| LLaVA |
-0 |
-286 |
-98 |
-3 |
-13 |
-54.4 |
-52.3 |
-331 |
-134 |
-13 |
-2 |
-69.7 |
-67.7 |
-363 |
-112 |
-28 |
-1 |
-61.5 |
-58.1 |
-116 |
-65 |
-2 |
-3 |
-56.3 |
-55.9 |
-1 |
-197 |
-163 |
-7 |
-33 |
-48.6 |
-48.5 |
-169 |
-236 |
-15 |
-60 |
-69.6 |
-66.3 |
-247 |
-211 |
-9 |
-37 |
-57.6 |
-57.9 |
-68 |
-105 |
-5 |
-8 |
-56.6 |
-57.5 |
-2 |
-170 |
-223 |
-4 |
-3 |
-49.1 |
-49.5 |
-197 |
-274 |
-7 |
-2 |
-69.0 |
-68.5 |
-318 |
-178 |
-7 |
-1 |
-58.1 |
-57.9 |
-41 |
-141 |
-4 |
-0 |
-56.0 |
-57.0 |
-3 |
-243 |
-145 |
-8 |
-4 |
-52.6 |
-53.0 |
-243 |
-222 |
-13 |
-2 |
-71.8 |
-70.2 |
-295 |
-197 |
-11 |
-1 |
-57.3 |
-56.3 |
-82 |
-93 |
-10 |
-1 |
-56.5 |
-56.5 |
-4 |
-211 |
-180 |
-6 |
-3 |
-62.1 |
-62.0 |
-232 |
-244 |
-3 |
-1 |
-77.3 |
-77.1 |
-210 |
-286 |
-4 |
-4 |
-57.3 |
-56.3 |
-55 |
-129 |
-2 |
-0 |
-57.6 |
-57.0 |
-5 |
-241 |
-154 |
-3 |
-2 |
-63.3 |
-63.0 |
-313 |
-163 |
-4 |
-0 |
-80.5 |
-80.4 |
-260 |
-236 |
-8 |
-0 |
-62.1 |
-62.3 |
-109 |
-75 |
-2 |
-0 |
-52.7 |
-52.7 |
-4.3 RQ3: Robustness to Prompts
-Multilingual prompts
-Since our datasets primarily originate in English-language websites, multilingual prompts can assess the adaptability of intrinsic knowledge across different languages and mitigate data leakage.
-We translate our prompt template and claims into Chinese, German, and Arabic using the Google Translate API.
-Table 3 presents the performance across different languages.
-Notable variations can be observed in adherence to instructions, particularly in Arabic, which exhibits a higher frequency of Deny responses across datasets.
-Nevertheless, the overall accuracy levels remain comparable, suggesting that the responses’ lack of explanations and confidence scores in Arabic did not significantly impact model performance.
-Moreover, the True/False Accuracy generally surpasses the Overall Accuracy, indicating GPT-4V’s good knowledge awareness.
-We do not conduct experiments on open-source models as they are predominantly trained and tested in English.
-Prompts with similar semantics
-We employ ChatGPT to generate semantically similar yet distinct prompts, as depicted in Figure 3.
-The part pertaining to explanations and confidence scores is identical to the original prompt.
-We randomly select two prompts for testing GPT-4V and all generated prompts for the LLaVA(13b) model.
-The outcomes of the experiments are detailed in Table 4, where Index 0 denotes the original prompt, and the other indices represent the newly crafted prompts.
-The data reveals that GPT-4V maintains commendable consistency, while LLaVA demonstrates variable biases across different prompts.
-For the same dataset, such as Fauxtography, models tend to predict the majority of claims as True when using the original prompt (index 0).
-In contrast, the prompt of index 2 leads the models to predict most claims as False.
-Moreover, on Fauxtography and COSMOS, the maximum discrepancy in overall accuracy reached and , respectively, compared to just and for GPT-4V.
-It indicates that the LLaVA(13b) model is highly sensitive to prompts and fails to maintain consistency.
-4.4 RQ4: Improving Capabilities via In-Context Learning
-| Fauxtography | COSMOS | MOCHEG | Post-4V | ||||||||||||||||||||||
-Demo |
-Num |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-T |
-F |
-U |
-D |
-T/F |
-O |
-0 |
-0 |
-286 |
-98 |
-16 |
-0 |
-54.4 |
-52.8 |
-331 |
-134 |
-14 |
-1 |
-69.7 |
-67.7 |
-363 |
-112 |
-29 |
-0 |
-61.5 |
-58.1 |
-116 |
-65 |
-5 |
-0 |
-56.4 |
-55.9 |
-| 1 |
-1 |
-244 |
-135 |
-9 |
-12 |
-60.9 |
-61.8 |
-295 |
-171 |
-7 |
-7 |
-75.3 |
-74.2 |
-255 |
-229 |
-12 |
-8 |
-62.6 |
-62.7 |
-95 |
-88 |
-3 |
-0 |
-56.3 |
-55.4 |
-2 |
-182 |
-161 |
-6 |
-51 |
-59.5 |
-60.0 |
-234 |
-220 |
-8 |
-18 |
-75.8 |
-73.5 |
-169 |
-282 |
-10 |
-43 |
-57.4 |
-55.5 |
-42 |
-109 |
-6 |
-29 |
-67.5 |
-63.4 |
-| 2 |
-1 |
-264 |
-121 |
-4 |
-11 |
-61.6 |
-62.8 |
-337 |
-131 |
-7 |
-5 |
-73.9 |
-73.5 |
-314 |
-180 |
-4 |
-6 |
-64.8 |
-65.1 |
-97 |
-79 |
-10 |
-0 |
-57.4 |
-55.9 |
-2 |
-176 |
-204 |
-10 |
-10 |
-60.0 |
-60.5 |
-233 |
-230 |
-11 |
-6 |
-80.3 |
-79.6 |
-199 |
-291 |
-8 |
-6 |
-61.2 |
-60.7 |
-44 |
-131 |
-11 |
-0 |
-60.0 |
-60.8 |
-| 3 |
-1 |
-211 |
-185 |
-1 |
-3 |
-62.6 |
-62.8 |
-279 |
-195 |
-6 |
-0 |
-72.2 |
-71.7 |
-236 |
-266 |
-2 |
-0 |
-61.8 |
-61.5 |
-75 |
-111 |
-0 |
-0 |
-60.2 |
-60.2 |
-2 |
-191 |
-198 |
-9 |
-2 |
-64.1 |
-63.8 |
-263 |
-201 |
-16 |
-0 |
-75.4 |
-75.4 |
-206 |
-288 |
-9 |
-1 |
-61.1 |
-61.7 |
-54 |
-127 |
-4 |
-1 |
-59.7 |
-59.7 |
-| 4 |
-1 |
-191 |
-193 |
-9 |
-7 |
-62.2 |
-62.0 |
-260 |
-203 |
-16 |
-1 |
-78.0 |
-77.9 |
-210 |
-283 |
-11 |
-0 |
-60.6 |
-60.9 |
-50 |
-126 |
-8 |
-2 |
-60.8 |
-60.8 |
-2 |
-196 |
-188 |
-7 |
-9 |
-62.8 |
-62.8 |
-259 |
-207 |
-12 |
-2 |
-77.5 |
-77.5 |
-207 |
-284 |
-11 |
-2 |
-61.1 |
-61.3 |
-52 |
-122 |
-10 |
-2 |
-61.5 |
-61.8 |
-As demonstrated in the 1st and 4th examples in Figure 2, the LLaVA(13b) model struggles to distinguish claims associated with manipulated images.
-The image content is consistent with the textual claim but contradicts real-world facts.
-To enhance the model’s comprehension of the task, checking facts rather than the consistency between the image and text, we experiment with employing in-context learning to improve its performance.
-In detail, we obtain GPT-4V’s responses for claims outside our test datasets and employ these as instructive examples for LLaVA(13b).
-Due to LLaVA’s limitation with multiple image inputs, we combine different images vertically with a 20-pixel gap between them.
-The linkage between images and claims is narratively indicated, as illustrated in the following prompt:
+| T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | |
+| GPT-4V | 158 | 195 | 29 | 18 | 81.9 | 82.1 | 179 | 204 | 83 | 14 | 86.2 | 80.0 | 216 | 223 | 37 | 28 | 87.2 | 83.5 | 54 | 98 | 26 | 8 | 79.6 | 73.7 |
+| MiniGPT | - | - | - | - | - | 55.5 | - | - | - | - | - | 62.1 | - | - | - | - | - | 54.3 | - | - | - | - | - | 55.9 |
+| LLaVA(7b) | 337 | 6 | 1 | 56 | 53.4 | 46.3 | 449 | 7 | 4 | 20 | 52.0 | 50.0 | 409 | 6 | 0 | 89 | 54.2 | 44.8 | 157 | 1 | 3 | 25 | 42.0 | 37.1 |
+| LLaVA(13b) | 286 | 98 | 3 | 13 | 54.4 | 52.3 | 331 | 134 | 13 | 2 | 69.7 | 67.7 | 363 | 112 | 28 | 1 | 61.5 | 58.1 | 116 | 65 | 2 | 3 | 56.4 | 55.9 |
+| LLaVA+PE | 244 | 153 | 2 | 1 | 57.1 | 54.7 | 275 | 204 | 0 | 1 | 76.3 | 71.7 | 290 | 214 | 0 | 0 | 59.9 | 58.1 | 85 | 101 | 0 | 0 | 56.9 | 56.1 |
+| LLaVA+ICL-1 | 228 | 159 | 6 | 8 | 61.8 | 62.3 | 293 | 175 | 9 | 3 | 74.8 | 74.1 | 254 | 240 | 7 | 4 | 62.4 | 62.5 | 79 | 101 | 5 | 1 | 58.7 | 57.5 |
+| LLaVA+ICL-2 | 186 | 188 | 8 | 18 | 61.6 | 61.7 | 247 | 215 | 12 | 7 | 77.3 | 76.5 | 195 | 286 | 10 | 13 | 60.2 | 60.4 | 48 | 122 | 8 | 8 | 62.1 | 61.4 |
+| Majority | - | - | - | - | - | 50.0 | - | - | - | - | 50.0 | - | - | - | - | - | 53.0 | - | - | - | - | - | 56.5 | |
+3.3 Evaluation Metrics
+Response Types
+Based on the three components (prediction, explanation, and confidence level) in the model’s response, we categorize them into four types.
+Others refer to instances where the model fails to respond according to our requirements, missing any component.
+Uncertain denotes cases where the model explicitly expresses uncertainty, with statements such as “cannot confirm”, “cannot verify”, or “cannot determine”, etc.
+The remaining samples are grouped as True or False according to the model’s prediction.
+Overall, the numbers of these four indicators reveal the model’s ability to follow instructions, express uncertainty, and inherent bias in predicting True and False.
+Accuracy Metrics
+We design two accuracy metrics to reflect the model’s performance.
+One is True&False Accuracy, measuring the accuracy of samples that are solely classified as True or False.
+This is important because users usually do not rely on the model’s response when the model clearly expresses uncertainty or does not follow instructions.
+The other metric, named Overall Accuracy, calculates the proportion of responses that contain correct predictions across all samples.
+This facilitates comparison between different approaches.
+4 Experimental Setups
+We use OpenAI’s API to collect responses from GPT-4V (gpt-4-1106-preview).
+In addition, we experiment with open-source MLLMs, including LLaVA-1.5 (7b and 13b, Liu et al.
+2023b, a), and MiniGPT (v2, Chen et al.
+2023).
+All parameters are set to the default values, with max_tokens=300 for GPT-4V.
+These models are notably representative and competitive in a multimodal setup Fu et al.
+(2023); Liu et al.
+(2023c).
+We further explore two approaches that do not require fine-tuning to enhancing the model’s performance:
+Prompt Ensembles (PE)
+involve using a variety of prompts for the same task and aggregating model’s responses to produce more accurate and less biased predictions.
+We employ ChatGPT to generate 5 more semantically similar yet distinct prompts, as depicted in Figure 2.
+The part about explanations and confidence scores is identical to the original prompt.
+With these prompts we collect six responses for each claim.
+Then, we conduct a vote to determine the final response type.
+If the two highest-scoring response types receive the same number of votes, we consider the model uncertain.
+In-Context Learning (ICL)
+enables a model to learn tasks by directly observing examples within the prompt, bypassing the need for prior explicit training.
+Debunking articles, which are lengthy and often contain irrelevant descriptions, are not used here.
+Instead, we utilize responses from GPT-4V to claims not included in our test datasets as instructive examples.
+To overcome LLaVA’s limitation in handling multiple image inputs, we vertically stack different images with a 20-pixel separation.
+The relationship between each image and its corresponding claim is explicitly described, such as "For the first picture, claim: ", as shown in the prompt below.
+We perform 1 and 2-shot learning, showing the model one and two examples, respectively, referred to as ICL-1 and ICL-2.
+We use four pairs of examples as demonstrations.
+For ICL-1, all examples are False, while for ICL-2, the first claim is always False, and the second is True.
+They are detailed illustrated in Figure 10 in Appendix.
+0.5cm0.5cm
 Use the following format to answer whether the claim regarding the pictures is True or False:
@@
 For the third picture, claim: "CLAIM"
-We construct four sets of demonstrations, each comprising one true and one false claim.
-We report the performances when using only the false claim and employing both claims as examples.
-We observe a universal improvement in model accuracy when examples are presented in Table 5.
-However, the use of the second example does not contribute to accuracy.
-In the explanations provided by LLaVA, we can observe two effects: one is the average length of the explanation.
-For instance, on MOCHEG, with the presentation of 0, 1, and 2 examples from the first set, the average sentence length of the model’s explanations are 3.09, 4.02, and 3.75, respectively.
-Another change is the frequency of the term "manipulated," which increases from 4 to 170 and 193 on MOCHEG.
-This suggests that the model can learn to analyze from the perspective of image manipulation.
-The last two rows of Figure 2 display responses from the LLaVA(13b) model.
-In the 4th example, it is evident that through learning, the model recognizes the altered part in the image and explains the motive behind the modifications: "but the gun and menacing pose are added elements," and "image is likely created to spread misinformation and discredit Joe Biden or to provoke a reaction from viewers." However, some samples are not correctly identified, such as the 1st example, and we observe that LLaVA produces outdated information, as highlighted in green.
-4.5 RQ5: Calibration Evaluation
-Calibration refers to the alignment between a model’s predicted probabilities (or confidence levels) and the actual outcomes.
-It reflects the model’s knowledge-awareness and ability to accurately estimate the confidence in its responses Geng et al.
-(2023).
-In this evaluation, we use verbalized confidence, which is readily accessible.
-Figure 4 displays the confidence distributions alongside calibration curves for GPT-4V and LLaVA(13b), breaking down the data into 10% confidence intervals.
+5 Experimental Results
+We comprehensively analyze the data we collected in Section 5.1.
+In Section 5.2, we provide a detailed analysis of the model explanations, focusing on their step-by-step reasoning, and discuss our taxonomy of failure reasons.
+Section 5.3 is dedicated to ablation studies.
+5.1 Comprehensive Results
+The comprehensive results are displayed in Table 2.
+The first four columns for each dataset represent the number of different types of responses, while the following are True&False and Overall Accuracy, respectively.
+We have highlighted the highest metrics and the highest ones among the open-source models.
+Hereafter, we omit the version numbers of the open-source models and refer to them simply as LLaVA(7b), LLaVA(13b), and MiniGPT.
+Both PE and ICL experiments are based on LLaVA(13b).
+LLaVA+PE involves voting based on 6 responses; if the counts of True and False are equal, then it is an uncertain case.
+LLaVA+ICL-1 and LLaVA+ICL-2 represent the average results across four sets of demonstrations, we have rounded the number of different responses.
+Detailed outcomes for PE and ICL are presented in Table 5 and 6 in Appendix.
+Overall, GPT-4V demonstrates surprising accuracy, with both accuracy metrics exceeding 80% on the three public datasets.
+The accuracy decrease is not particularly significant on Post-4V, with the True&False Accuracy equals .
+In addition, these values are mostly higher than Overall Accuracy, suggesting that GPT-4V exhibits higher precision when it responds with confidence.
+There are more False cases than True ones.
+The high number of Others is primarily due to GPT-4V safety alignment.
+This results in frequent replies such as "I’m sorry, but I cannot assist with this request." or "This content may violate our content policy." Additionally, GPT-4V expresses uncertainty the most.
+In contrast, MiniGPT cannot provide explanations and confidence, while LLaVA(7b) shows a strong bias, almost always responding True with this prompt.
+LLaVA(13b) shows improved accuracy across all datasets, better adherence to instructions, and more frequent expressions of uncertainty.
+We observe that the two approaches, PE and ICL, can enhance the model’s performance to varying degrees.
+As to PE, there is an obvious improvement in the COSMOS dataset, with increases of 6.6 and 4.0 in two accuracy metrics, respectively.
+However, the enhancement on other datasets is very limited.
+ICL shows a greater impact.
+On the COSMOS dataset, LLaVA+ICL-1 brings increases of 7.6 and 9.0 in both metrics.
+On MOCHEG, the improvement is minimal, with gains of 0.9 and 4.4, respectively.
+Calibration
+The left part of Figure 3 displays the confidence distributions of GPT-4V and LLaVA(13b), breaking down the data into 10% confidence intervals.
 For instance, a value of 95 corresponds to confidence levels within the bracket.
 GPT-4V’s confidence scores are largely clustered in the interval, whereas LLaVA(13b)’s scores are more frequently found in the range.
-The calibration curves reveal that both models exhibit a positive correlation between confidence levels and accuracy, with GPT-4V’s calibration closely mirroring the ideal curve (illustrated by the dashed line), indicating well-calibrated confidence scores.
+The right calibration curves reveal that both GPT-4V and LLaVA exhibit a positive correlation between confidence levels and accuracy, with GPT-4V’s calibration closely mirroring the ideal curve (illustrated by the dashed line), indicating well-calibrated confidence scores.
 Conversely, LLaVA(13b)’s curve suggests a propensity for overconfidence.
-The Expected Calibration Errors (ECEs, Guo et al.
-2017), detailed in Table 6, further substantiate that GPT-4V maintains superior calibration.
-ECE |
-Fauxtography |
-COSMOS |
-MOCHEG |
-Post-4V |
-GPT-4V |
-7.76 |
-4.36 |
-4.88 |
-11.83 |
-LLaVA |
-30.20 |
-15.26 |
-22.82 |
-28.45 |
-4.6 RQ6: Taxonomy of Failure Reasons
+We find that ICL does not result in better calibration of LLaVA’s verbalized confidence as their curves are very close.
+5.2 Explanation-based Analysis
+5.2.1 Case Studies
+Figure 4 displays totally 4 cases with the original multimodal claims and corresponding responses .
+Rows labeled "GPT-4V w.
+Image," "LLaVA w.
+Image," and "LLaVA w.
+Image 1 example" correspond to the results of GPT-4V, LLaVA(13b), and LLaVA+ICL-1, respectively.
+It can be observed that GPT-4V’s explanation contains a richer set of facts.
+For instance, it explicitly states, "there was never an official campaign for Michelle Obama in the 2020 election," in the 1st case, and in the 4th case, it provides detailed background information about Margaret Hamilton, including her role in "leading the development of onboard flight software for the Apollo mission." Moreover, GPT-4V adeptly incorporates details from the images, such as referencing a "Google image search: Creepy Joe Biden" in the 3rd image.
+However, LLaVA(13b) lacks the capability to counteract manipulated images, as evidenced in the 1st and 3rd cases.
+Additionally, LLaVA(13b)’s explanations are typically brief, often limited to a description of the image followed by a conclusion about its veracity.
+It shows a lack of effective reasoning, rendering the explanations less comprehensive than GPT-4V.
+We find ICL enhances LLaVA(13b)’s fact-checking capabilities.
+For instance, the model accurately analyzes the 3rd image, with "a genuine photograph of Joe Biden and an elderly woman, but the gun and menacing pose are added elements that are not part of the original scene." In the 4th case, LLaVA+ICL-1 provides a more detailed description of Margaret Hamilton’s background and states that "the image is an authentic representation of her contribution." However, it still fails in the 1st case with ICL.
+5.2.2 Analysis of Explanation Length
+For convenience, we refer to the number of sentences in the explanation part of a response as the explanation length.
+The average length of model explanations with different approaches is shown in Figure 5.
+We use NLTK to segment the sentences.
+GPT-4V provides the highest number of sentences in explanations across various datasets, with a minimum of 3.57 sentences in COSMOS and a maximum of 3.88 in Fauxtography.
+The value of LLaVA+PE represents the average length of explanations under different prompts.
+ICL prompts LLaVA to generate more sentences; however, providing two examples (LLaVA+ICL-2) does not yield longer explanations than providing just one (LLaVA+ICL-1), the average length even decreases.
+We further analyze the relationship between the model’s prediction accuracy and the explanation length, as depicted in Figure 6, where we display explanation lengths ranging from 2 to 7, encompassing the most samples generated by various models.
+Typically, different models achieve peak accuracy when the explanation length is 5 or 6 sentences.
+Specifically, on COSMOS, the accuracy of the different models grows steadily as the explanation length increases from 2 to 5.
+Across the other three datasets, the left half of the curves all show varying degrees of ascent despite noticeable fluctuations.
+GPT-4V achieves its highest accuracy at an explanation length of 5 sentences on them.
+| Fauxtography | COSMOS | MOCHEG | Post-4V | |||||||||||||||||||||
+| T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | |
+| GPT-4V | 105 | 202 | 32 | 61 | 80.1 | 71.0 | 111 | 253 | 50 | 66 | 64.0 | 57.5 | 153 | 265 | 38 | 48 | 73.9 | 68.9 | 35 | 68 | 10 | 73 | 74.8 | 67.7 |
+| LLaVA(7b) | 398 | 0 | 2 | 0 | 50.3 | 50.0 | 474 | 0 | 4 | 2 | 50.2 | 49.8 | 502 | 0 | 2 | 0 | 53.0 | 52.8 | 181 | 0 | 4 | 1 | 43.1 | 43.0 |
+| LLaVA(13b) | 167 | 214 | 17 | 2 | 62.5 | 62.0 | 263 | 195 | 20 | 2 | 76.6 | 75.8 | 168 | 316 | 20 | 0 | 62.0 | 61.5 | 40 | 124 | 22 | 0 | 60.4 | 60.8 |
+5.2.3 Reasons for Predicting False
+To better understand the model’s fact-checking process, particularly why it labels multimodal claims as false, we explore two types of cases: checked cases, the model confirms that statements have been verified by organizations; manipulated cases, it indicates that images have been altered.
+Checked Cases
+We count the occurrences of phrases such as "been checked" and "been debunked" across different settings, with the results displayed in left graph in Figure 7.
+GPT-4V mentions these keywords most frequently across three public datasets, significantly more than others.
+Additionally, ICL prompts LLaVA to mention them more often.
+Notably, our ICL examples do not include checked cases.
+However, on Post-4V, several claims are considered fact-checked by LLaVA, but not by GPT-4V.
+The accuracy of GPT-4V and LLaVA+ICL on checked cases is and , respectively.
+The average True&False Accuracy for those that are not checked cases is only 84.4% and 65.7%, respectively.
+Manipulated Cases
+We notice that only the examples in the 3rd demonstration do not include manipulated cases.
+Therefore, we compare the impact of demonstrations on the number of manipulated cases, as shown in the right of Figure 7.
+Despite not accessing manipulated cases as example, the number significantly rises in demo3-1-exp1 and demo3-exp2, although these figures remain lower compared to other demonstrations.
+Thus, reasoning from the perspective of image manipulation is also not an explicit learning process.
+The accuracy of LLaVA is 67.6% for manipulated cases and 65.4% for non-manipulated cases, respectively.
+This demonstrates LLaVA’s potential to analyze veracity from the perspective of image manipulation.
+| Fauxtography | COSMOS | MOCHEG | Post-4V | |||||||||||||||||||||
+| T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | T | F | U | O | T&F | All | |
+| GPT-4V | 158 | 158 | 72 | 12 | 80.7 | 63.8 | 162 | 190 | 109 | 18 | 79.8 | 70.0 | 234 | 204 | 54 | 33 | 83.7 | 77.8 | 51 | 83 | 38 | 14 | 80.1 | 72.6 |
+| LLaVA(7b) | 237 | 71 | 7 | 49 | 54.1 | 50.5 | 311 | 101 | 14 | 54 | 66.0 | 60.1 | 341 | 89 | 8 | 66 | 56.0 | 50.8 | 125 | 37 | 3 | 21 | 51.2 | 49.5 |
+| LLaVA(13b) | 207 | 143 | 7 | 42 | 51.7 | 52.0 | 228 | 201 | 22 | 28 | 69.9 | 68.5 | 235 | 212 | 7 | 50 | 61.8 | 59.5 | 65 | 105 | 2 | 14 | 61.8 | 60.2 |
+5.2.4 Taxonomy of Failure Reasons
 Upon examining GPT-4V’s explanations, we identify six primary categories of failure reasons.
-Failures attributed to Lack of Knowledge occur when responses are based on trivial or incomplete information, often contradicting common sense.
-Recognition Errors arise from the model’s visual module misinterpreting images, including failing to identify altered or AI-generated visuals.
-Reasoning Errors occur when the model incorrectly interprets images due to factors like unnatural distortions or unusual perspectives, or there are self-contradictions in reasoning.
-Failures due to Partial Facts happen when the model addresses only a subset of the facts present in a textual claim.
-Image Issues involve failures related to poor image quality, such as blurriness or incompleteness, affecting the model’s ability to recognize content accurately.
+Lack of Knowledge, especially obscure knowledge that contradicts intuition, easily leads the model to assess a statement’s truthfulness inaccurately.
+Similar Images refer to instances where the model associates with another image featuring a similar scene.
+In such cases, the model incorrectly extracts knowledge, leading to inconsistent content with the presented image.
+Distortion Due to Perspective and Lens, we find that GPT-4 cannot correctly understand images that are distorted due to the shooting angle or lens..
+Incomplete Verification refers to the cases where the model does not fully consider the atomic claims within a multimodal claim.
+For example, when a claim contains both correct and incorrect subclaims, or when the textual claim is accurate but the accompanying image is misleading.
+Blurry or Incomplete Image can lead to an incorrect prediction due to the model’s uncertainty about what it sees.
 Ambiguity Claims involve semantic uncertainties, where a term could reference multiple entities.
-Examples beyond the lack of knowledge are presented in Figure 5.
-In subfigure (a), GPT-4V misinterprets scars from a dog’s burn as being painted or digitally altered.
-In subfigure (b), it correctly identifies the locations of Notre Dame and the Eiffel Tower but engages in erroneous reasoning.
-In subfigure (c), the model understands that a kiss can release hormones like oxytocin and knows that MR images cannot capture chemical reactions, with its predictions primarily influenced by the first fact.
-In subfigure (d), the model cannot discern that the scene depicted is related to posture training based on the image’s content.
-In subfigure (e), "Playboy" has multiple meanings, which could refer to a publisher or magazine’s name.
-We also manually tall the frequency of these errors across various datasets, as detailed in Table 7.
-Overall, the primary cause of errors is attributed to Lack of Knowledge, suggesting that augmenting the system with additional evidence could enhance its performance.
-There is room for improving the model’s ability to understand and reason about AI-generated or distorted images.
-Know |
-Recog |
-Reas |
-Partial |
-Image |
-Ambi |
-| Fauxtography |
-59.3 |
-20.3 |
-17.2 |
-1.6 |
-1.6 |
-0.0 |
-| COSMOS |
-80.3 |
-5.9 |
-9.8 |
-2.0 |
-0 |
-2.0 |
-| MOCHEG |
-73.2 |
-3.6 |
-7.1 |
-3.6 |
-5.4 |
-7.1 |
-| Post-4V |
-85.2 |
-11.1 |
-0.0 |
-0.0 |
-0.0 |
-3.7 |
-We also explore the utilization of GPT-4 for detecting nonfactual content within explanations when fact-checking articles are available.
-We design a prompt comprising two steps: removing sentences containing no facts and predicting whether the reference article supports the sentence.
-Due to the limited content space of the article, we do not present the specific prompt.
-Figure 6 showcases two instances where, although the predictions are correct, the facts within the explanations are inaccurate.
-This underscores the necessity for further examination of their veracity.
-The photograph on the left, featuring a young girl named Fouzia, was captured in Morocco by Giacomo Pirozzi, a photographer for the United Nations Children’s Fund (UNICEF).
-However, GPT-4V incorrectly attributes it to photographer Terri Gold in Yemen, 1998.
-The right photo is incorrectly identified as from the movie "Platoon," actually, it is from the movie "Tropic Thunder." The returned sentences also contained substantial image descriptions not present in the original fact-checking articles.
-Improving the automatic assessment of the factuality in explanations is a focus for our future work.
+We show representative examples and explanations for each failure reason in Figure 8.
+Our taxonomy can lead to targeted optimization strategies.
+For instance, we could include more images with distortions or blurry or incomplete images from image scaling and cropping into the instruction-tuning dataset.
+5.3 Ablation Studies
+Impact of Images
+To understand the impact of images on fact-checking, we conduct a comparative experiment where we ask GPT-4V and LLaVA (13b) to re-evaluate the veracity of textual claims without providing images.
+As shown in Table 3, the results are marked in red for values that exceed those in Table 2.
+GPT-4V experiences a decrease in accuracy across all datasets, while surprisingly, LLaVA (13b) shows higher accuracy compared to when images are included.
+To understand this, we further look into their explanations.
+As shown in Figure 4, in the 1st and 3rd cases, the model can make correct predictions based on internal knowledge without images, stating "it is unlikely that Barack Obama would be endorsing his wife’s …" and "Joe Biden is a public figure and it is highly unlikely that he would pose menacingly …" However, the model is misled when manipulated images are presented.
+The GPT-4V’s responses without image input are in Figure 9 in Appendix.
+Reasoning First
+In the experiments above, we prompt the model to first output a prediction, followed by step-by-step reasoning.
+We conduct a comparative experiment to assess the impact of reversing this order—starting with reasoning before prediction—on accuracy.
+The results are presented in Table 4, where values exceeding those in Table 2 are highlighted in red.
+We observe that starting with reasoning significantly increases the instances of GPT-4V expressing uncertainty.
+For example, the number of uncertain cases rises from 29 to 72 on Fauxtography and 10 to 38 on Post-4V.
+This increase directly leads to an obvious decline in Overall Accuracy, although True&False Accuracy remains relatively high.
+LLaVA(7b) still predicts most claims as True.
+LLaVA(13b) shows varying accuracy improvements across the last three datasets, though the changes are not substantial.
+We also note an increase in instances where the model fails to follow instructions, which may partially influence the accuracy metrics.
+6 Limitations
+We acknowledge the following limitations in our work.
+First, due to constraints in manpower and computational resources, we did not test the complete Fauxtography and COSMOS datasets.
+The limited number of samples may impede an accurate assessment of model characteristics.
+Second, we did not quantify the different capabilities within fact-checking, such as the ability of different models to recognize AI-generated images, identify faces and scenes, and retrieve contextual knowledge.
+Moreover, we did not investigate how fine-tuning with domain-specific data could potentially improve model performance.
+Future research will aim to address these gaps and provide a more comprehensive evaluation of the models.
 7 Conclusion and Future Work
@@
 Our results indicated that state-of-the-art MLLMs, such as GPT-4V, have the potential to assist professional fact-checkers: they can enhance their efficiency by providing reference predictions, valuable clues, and explanatory insights, together with confidence.
-Open-source models exhibit sensitivity to text, with different prompts revealing various biases.
-Moreover, they are susceptible to being deceived by altered images, and even in the absence of images, they tend to produce hallucinations pertaining to image descriptions.
-Our work demonstrates both the potential and limitations of MLLMs, laying the groundwork for further leveraging MLLMs to support fact-checking.
+However, open-source models like LLaVA fall behind.
+They can be misled by manipulated images and may generate hallucinations when no image is present.
+We empirically demonstrate that prompt ensembles (PE) and in-context learning (ICL) are able to improve model’s accuracy in detecting misinformation, while ICL is more efficient.
 In future work, we plan to investigate how to enhance the fact-checking capabilities of models, focusing not only on improving accuracy but also on strengthening robustness and augmenting the knowledge reasoning and the verification abilities required for fact-checking.
-We further plan to explore how MLLMs can better support fact-checking when connected to external knowledge.
+A potential approach is to use GPT-4V to enhance the capabilities of smaller open-source models through knowledge distillation.
+We also plan to explore how MLLMs can better support fact-checking when connected to external knowledge.
+Appendix A More Experimental Results
+Figure 9 displays additional responses, where "w.
+Image" indicates multimodal claims as input and "w/o.
+Image" denotes text-only claims as input.
+The last two rows show the results for LLaVA+ICL-1 and LLaVA+ICL-2, respectively.
+As shown in Figure 9, GPT-4V’s explanations demonstrate its awareness of missing images, employing phrases such as "there is no verified public record" or "without any provided context." This indicates that GPT-4V actively retrieves knowledge to verify the presence or context of images.
+Compared with no demonstrations, the results reveal that ICL enhances the model’s reasoning capabilities, incorporating more detailed analysis of images and discussions about the motivations behind their creation.
+Table 5 shows the results with different prompts in prompt ensembles (PE).
+The accuracy of the GPT-4V model is quite consistent across various prompts.
+In contrast, the LLaVA model is very sensitive to the prompts, with some prompts showing significantly higher accuracy than others on the same dataset.
+However, no single prompt consistently performs best across all datasets.
+Table 6 displays the impact of different demonstrations in ICL on the LLaVA (13b) model.
+Demonstrations clearly enhance the model’s accuracy.
+However, the benefit of the second example is limited, and it does not show a significant improvement compared to the ICL-I cases.